OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (anonymous):

use the rule: y2-y1=m(x2-x1)

OpenStudy (anonymous):

Could you explain more?

OpenStudy (anonymous):

(1, 5) and (-2 , 3) ^ ^ ^ ^ x1 ,y1 x2 y2

OpenStudy (anonymous):

Got it ?

OpenStudy (anonymous):

ax+by=c is the standard form. So do as told^^ and then write your answer in the form I wrote.

OpenStudy (anonymous):

That is up to the learner...If they can't do it then I will assist further. I just don't give out answers.

OpenStudy (anonymous):

Ok, @crisscross, find the slope of the line first using the given points and the equation \[(y_2-y_1)/(x_2-x_1)\] and then pick any (x,y) point from the two that you have and do \[(y-5)= slope \times (x-1)\] This will give you your answer in the form of y=mx+c. So just move mx to the left hand side of the equation to have it in standard form.

OpenStudy (anonymous):

@Eyad Math never gives me trouble mate.

OpenStudy (anonymous):

I dont think you are getting my point here. STANDARD FORM is what they want. y=mx+c is not the standard form.

OpenStudy (anonymous):

No it is not, that's in slope intercept form